OpenStudy (tanya123):

How can I get the answer 24 by only using the numbers 8,8,3,3. You can use the main signs add, subtract multiply and divide.

OpenStudy (zzr0ck3r):

do you have to use them all and can you use them ore than once?

OpenStudy (zzr0ck3r):

ie 8+8+8 is obviously not what we ware looking for

OpenStudy (kainui):

Well 8*3 is 24, throw away the other 8 and 3 and you're done lol.

OpenStudy (nincompoop):

8*3 = 8*3

OpenStudy (anonymous):

that's a silly question @zzr0ck3r, the numbers are repeating in the list

OpenStudy (zzr0ck3r):

ok @Abmon98 thnx

OpenStudy (astrophysics):

Oh I've seen this, but you could use logarithms and factorials, which made it much easier.

OpenStudy (anonymous):

take the square root if you have to use them all

OpenStudy (kainui):

Well only +,-, *,/ right?

OpenStudy (zzr0ck3r):

\(\frac{8}{3-\frac{8}{3}}\)

OpenStudy (zzr0ck3r):

its always an answer like that^

OpenStudy (zzr0ck3r):

in fairness I have seen these a bunch

OpenStudy (astrophysics):

That's right zz ^.^

OpenStudy (anonymous):

got it 3/(3/8) * 3

OpenStudy (zzr0ck3r):

someone times they allow bonus rules and you can get it a different way, like square roots, factorials, ...

OpenStudy (zzr0ck3r):

some times*

OpenStudy (zzr0ck3r):

those are harder:)

OpenStudy (anonymous):

:P

OpenStudy (kainui):

Hmm interesting, this looks like one of the continued fraction representations of tangles in knot theory.

OpenStudy (nincompoop):

you have to use three twice and eight twice

OpenStudy (abmon98):

8/(3-(8/3)) = 8/(1/3)

OpenStudy (anonymous):

i was joking, nin but looks like abmon's got it, nice

OpenStudy (nincompoop):

8*3 = 8*3

OpenStudy (zzr0ck3r):

good job @nincompoop :P

OpenStudy (tanya123):

4)(8-3)!/(8-3) ( × )! ( + )! √(8×8×3×3) 8+(8×(3!/3)) ((√(8+8) × (3/3))! √(8+8) × (3+3) (log base(3!/3) of 8) × 8 ((log base(3!/3) of (8+8))

OpenStudy (zzr0ck3r):

up is up

OpenStudy (astrophysics):

Tanya forgot to mention, factorials and logarithms as I mentioned above XD

OpenStudy (kainui):

I like this one √(8+8) × (3+3)

OpenStudy (nincompoop):

(+)! is a nice emoticon (x).(+)

OpenStudy (anonymous):

https://answers.yahoo.com/question/index?qid=1006053013451 \[\frac{8}{3-\frac{8}{3}}=\frac{8}{\frac{1}{3}}=24 \]

OpenStudy (wearehiding):

good job.